The alternator on a truck engine produces 350Wof electric power when it rotates at 4000 rev/min. If it is 95 percent efficient, what should the difference (in newtons) between the tensions in the tight and slack parts of its V-belt drive be? The diameter of the alternator’s pulley is 10 cm. 


1
Expert's answer
2021-11-04T16:20:43-0400

Torque:


"\tau=\frac P\omega=5.25\text{ Nm}."

Tension:


"T=\frac \tau r=52.5\text{ N}."

The difference in tension:


"\Delta T=T(1-\eta)=2.63\text{ N}."